# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@ 6.2. Fourth moment bounds for $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)$
If $\vec{\mathrm{i}}$ appears for four times, then we must have
(7.19) $\begin{split}&{\mathcal{E}}_{W}^{4}(\vec{\mathrm{i}})=c_{1}^{4}{\mathcal{E}}_{\Theta}^{4}(\vec{\mathrm{i}})+6c_{1}^{2}c_{2}^{2}{\mathcal{E}}_{\Theta}^{2}(\vec{\mathrm{i}}){\mathcal{E}}_{G}^{2}(\vec{\mathrm{i}})+c_{2}^{4}{\mathcal{E}}_{G}^{4}(\vec{\mathrm{i}})\\ \leq C\big{(}c_{1}^{4}{\mathcal{E}}_{\Theta}^{4}((\mathrm{i},\mathrm{i}))&(g_{(\mathrm{i}^{\prime}-\mathrm{i})}(0))^{4}{\mathcal{E}}_{\Theta}^{4}((\mathrm{i}^{\prime},\mathrm{i}^{\prime}))+6c_{1}^{2}c_{2}^{2}{\mathcal{E}}_{\Theta}^{2}((\mathrm{i},\mathrm{i}))(g_{(\mathrm{i}^{\prime}-\mathrm{i})}(0))^{2}{\mathcal{E}}_{\Theta}^{2}((\mathrm{i}^{\prime},\mathrm{i}^{\prime})){\mathcal{E}}_{G}^{2}(\vec{\mathrm{i}})+c_{2}^{4}{\mathcal{E}}_{G}^{4}(\vec{\mathrm{i}})\big{)},\end{split}$
where we have used the same argument of step (A) in “adjustments on coarse-grained disorders and heat kernels” in the proof of Theorem 6.3 to replace $\Theta_{n}(\vec{\mathrm{i}})$ by $C\Theta_{n}((\mathrm{i},\mathrm{i}))g_{(\mathrm{i}^{\prime}-\mathrm{i})}(0)\Theta_{n}((\mathrm{i}^{\prime},\mathrm{i}^{\prime}))$.
Meanwhile, we have that
$\begin{split}{\mathbbm{E}}\Big{[}\Big{(}W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}((\mathrm{i},\mathrm{i}))g_{(\mathrm{i}^{\prime}-\mathrm{i})}(0)W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}((\mathrm{i}^{\prime},\mathrm{i}^{\prime}))\Big{)}^{4}\Big{]}={\mathcal{E}}_{W}^{4}((\mathrm{i},\mathrm{i}))(g_{(\mathrm{i}^{\prime}-\mathrm{i})}(0))^{4}{\mathcal{E}}_{W}^{4}((\mathrm{i}^{\prime},\mathrm{i}^{\prime})).\end{split}$
Note that by (7.18),
${\mathcal{E}}_{G}^{2}(\vec{\mathrm{i}})={\mathcal{E}}_{\Theta}^{2}(\vec{\mathrm{i}})\leq C{\mathcal{E}}_{\Theta}^{2}((\mathrm{i},\mathrm{i}))(g_{(\mathrm{i}^{\prime}-\mathrm{i})}(0))^{2}{\mathcal{E}}_{\Theta}^{2}((\mathrm{i}^{\prime},\mathrm{i}^{\prime}))=C{\mathcal{E}}_{G}^{2}((\mathrm{i},\mathrm{i}))(g_{(\mathrm{i}^{\prime}-\mathrm{i})}(0))^{2}{\mathcal{E}}_{G}^{2}((\mathrm{i}^{\prime},\mathrm{i}^{\prime})),$
and
${\mathcal{E}}_{G}^{4}(\vec{\mathrm{i}})=3({\mathcal{E}}_{G}^{2}(\vec{\mathrm{i}}))^{2}\leq C({\mathcal{E}}_{G}^{2}((\mathrm{i},\mathrm{i}))(g_{(\mathrm{i}^{\prime}-\mathrm{i})}(0))^{2}{\mathcal{E}}_{G}^{2}((\mathrm{i}^{\prime},\mathrm{i}^{\prime})))^{2}=C^{\prime}{\mathcal{E}}_{G}^{4}((\mathrm{i},\mathrm{i}))(g_{(\mathrm{i}^{\prime}-\mathrm{i})}(0))^{4}{\mathcal{E}}_{G}^{4}((\mathrm{i}^{\prime},\mathrm{i}^{\prime})).$
Hence, by (7.19),
${\mathbbm{E}}\Big{[}\Big{(}W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}})\Big{)}^{4}\Big{]}\leq C{\mathbbm{E}}\Big{[}\Big{(}W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}((\mathrm{i},\mathrm{i}))g_{(\mathrm{i}^{\prime}-\mathrm{i})}(0)W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}((\mathrm{i}^{\prime},\mathrm{i}^{\prime}))\Big{)}^{4}\Big{]},$
and we can replace $W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}})$ by $W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}((\mathrm{i},\mathrm{i}))g_{(\mathrm{i}^{\prime}-\mathrm{i})}(0)W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}((\mathrm{i}^{\prime},\mathrm{i}^{\prime}))$ as in the proof of Theorem 6.3.
If $\vec{\mathrm{i}}$ appears for three times, then at renewal time $\mathrm{i}$, there could be another renewal pair $\vec{\mathrm{i}}_{1}$ visits it, and at renewal time $\mathrm{i}^{\prime}$, there could also be another renewal pair $\vec{\mathrm{i}}_{2}$ visits it. Thus, we need to estimate
(7.20) $\begin{split}{\mathbbm{E}}\Big{[}\big{(}\Theta_{n}(\vec{\mathrm{i}}_{1})\big{)}^{k}\Big{(}W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}})\Big{)}^{3}\big{(}\Theta_{n}(\vec{\mathrm{i}}_{2})\big{)}^{l}\Big{]},\quad\text{for}~{}k,l=0,1.\end{split}$
The reason that we drop $\Theta_{n}^{\rm{G}}(\vec{\mathrm{i}}_{m})$ for $m=1,2$ is due to the fact that either they are absent or have zero means. The computation is similar to the above case and we only emphasize some key steps.
In the expansion of (7.20), we have the factorization
$\begin{split}&{\mathbbm{E}}\Big{[}\big{(}\Theta_{n}(\vec{\mathrm{i}}_{1})\big{)}^{k}\big{(}\Theta_{n}(\vec{\mathrm{i}})\big{)}^{3}\big{(}\Theta_{n}(\vec{\mathrm{i}}_{2})\big{)}^{l}\Big{]}\leq C{\mathbbm{E}}\Big{[}\big{(}\Theta_{n}(\vec{\mathrm{i}}_{1})\big{)}^{k}\big{(}\Theta_{n}((\mathrm{i},\mathrm{i}))\big{)}^{3}\Big{]}(g_{(\mathrm{i}^{\prime}-\mathrm{i})}(0))^{3}{\mathbbm{E}}\Big{[}\big{(}\Theta_{n}(\mathrm{i}^{\prime},\mathrm{i}^{\prime})\big{)}^{3}\big{(}\Theta_{n}(\vec{\mathrm{i}}_{2})\big{)}^{l}\Big{]},\end{split}$
and
@@
$\begin{split}&{\mathbbm{E}}\Big{[}\big{(}\Theta_{n}(\vec{\mathrm{i}}_{1})\big{)}^{k}\Theta_{n}(\vec{\mathrm{i}})\big{(}\Theta_{n}^{\rm(G)}(\vec{\mathrm{i}})\big{)}^{2}\big{(}\Theta_{n}(\vec{\mathrm{i}}_{2})\big{)}^{l}\Big{]}\\ \leq&C{\mathbbm{E}}\Big{[}\big{(}\Theta_{n}(\vec{\mathrm{i}}_{1})\big{)}^{k}\Theta_{n}((\mathrm{i},\mathrm{i}))\Big{]}(g_{(\mathrm{i}^{\prime}-\mathrm{i})}(0))^{3}{\mathbbm{E}}\Big{[}\Theta_{n}(\mathrm{i}^{\prime},\mathrm{i}^{\prime})\big{(}\Theta_{n}(\vec{\mathrm{i}}_{2})\big{)}^{l}\Big{]}{\mathcal{E}}_{\Theta}^{2}(\vec{\mathrm{i}})\\ \leq&C^{\prime}{\mathbbm{E}}\Big{[}\big{(}\Theta_{n}(\vec{\mathrm{i}}_{1})\big{)}^{k}\Theta_{n}((\mathrm{i},\mathrm{i}))\Big{]}(g_{(\mathrm{i}^{\prime}-\mathrm{i})}(0))^{5}{\mathbbm{E}}\Big{[}\Theta_{n}(\mathrm{i}^{\prime},\mathrm{i}^{\prime})\big{(}\Theta_{n}(\vec{\mathrm{i}}_{2})\big{)}^{l}\Big{]}{\mathcal{E}}_{\Theta}^{2}((\mathrm{i},\mathrm{i})){\mathcal{E}}_{\Theta}^{2}((\mathrm{i}^{\prime},\mathrm{i}^{\prime})).\end{split}$
All these terms in the expansion of (7.20) are contained in the expansion of
(7.21) ${\mathbbm{E}}\Big{[}\big{(}\Theta_{n}(\vec{\mathrm{i}}_{1})\big{)}^{k}\Big{(}W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}((\mathrm{i},\mathrm{i}))g_{(\mathrm{i}^{\prime}-\mathrm{i})}(0)W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}((\mathrm{i}^{\prime},\mathrm{i}^{\prime}))\Big{)}^{3}\big{(}\Theta_{n}(\vec{\mathrm{i}}_{2})\big{)}^{l}\Big{]},\quad\text{for}~{}k,l=0,1.$
If $\vec{\mathrm{i}}$ appears for two times, then similarly, we only need to show that all terms in the expansion of
(7.22) ${\mathbbm{E}}\Big{[}\big{(}\Theta_{n}(\vec{\mathrm{i}}_{1})\big{)}^{k}\Big{(}W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}})\Big{)}^{3}\big{(}\Theta_{n}(\vec{\mathrm{i}}_{2})\big{)}^{l}\Big{]}$
are contained in the expansion of
(7.23) ${\mathbbm{E}}\Big{[}\big{(}\Theta_{n}(\vec{\mathrm{i}}_{1})\big{)}^{k}\Big{(}W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}((\mathrm{i},\mathrm{i}))g_{(\mathrm{i}^{\prime}-\mathrm{i})}(0)W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}((\mathrm{i}^{\prime},\mathrm{i}^{\prime}))\Big{)}^{3}\big{(}\Theta_{n}(\vec{\mathrm{i}}_{2})\big{)}^{l}\Big{]},$
for $k,l=0,1,2$. Note that we drop the moments of $\Theta_{n}^{\rm(G)}(\vec{\mathrm{i}}_{m})$ for $m=1,2$ by independence.
@@
Finally, by the above replacement, the upper bound of (7.17) has the same structure of $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\Theta_{n})$. Proposition 6.5 also holds given that we have the same second and fourth moments bounds on $W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}})$. We are left to obtain analogues of Lemma 6.1 and Lemma 6.2.
For ${\mathcal{E}}_{W}^{4}(\vec{\mathrm{i}})$, we have (7.19). Then by Jensen’s inequality,
${\mathcal{E}}_{W}^{4}(\vec{\mathrm{i}})=c_{1}^{4}{\mathcal{E}}_{\Theta}^{4}(\vec{\mathrm{i}})+6c_{1}^{2}c_{2}^{2}({\mathcal{E}}_{\Theta}^{2}(\vec{\mathrm{i}}))^{2}+3c_{2}^{4}({\mathcal{E}}_{\Theta}^{2}(\vec{\mathrm{i}}))^{2}\leq(c_{1}^{4}+6c_{1}^{2}c_{2}^{2}+3c_{2}^{4}){\mathcal{E}}_{\Theta}^{4}(\vec{\mathrm{i}}),$
which provides a fourth moment bound similar to Lemma 6.2.
@@
Finally, since $c_{1},c_{2}$ can be $su\sqrt{t},u\sqrt{t}$ or $\sqrt{t}$ for $s,u,t\in[0,1]$, all constants here are bounded and the proof is completed. ∎
8. Proof for Theorem 1.10
For simplicity, we only consider the measure $u^{\delta}({\mathrm{d}}x):=\widetilde{u}_{0,1}^{\delta}({\mathrm{d}}x,\mathbbm{1}_{\mathbbm{R}})$. we write
(8.1) $u^{\delta}[f]:=\int_{\mathbbm{R}}f(x)u^{\delta}(1,x){\mathrm{d}}x.$
We show that $u^{\delta}[f]$ can be approximated by a coarse-grained model $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(f,\mathbbm{1}_{\mathbbm{R}})$ (see (2.46)) that will be specified later. Then the proof is completed by Section 7. To proceed, we will first introduce some preliminaries in the continuous setting and compute the second moment of $u^{\delta}[f]$.
8.1. Preliminaries
@@
(8.2) $\begin{split}u^{\delta}[f]=&\int_{\mathbbm{R}}f(x){\mathrm{d}}x+\sum\limits_{k=1}^{\infty}\beta_{\delta}^{k}\idotsint\limits_{0<t_{1}<\cdots<t_{k}<\delta^{-2}\atop x_{0},\cdots,x_{k}\in{\mathbbm{R}}}f(x_{0})g_{t_{1}}\big{(}x_{1}-\frac{x_{0}}{\delta}\big{)}\rho(x_{1}){\mathrm{d}}x_{0}{\mathrm{d}}x_{1}{\mathrm{d}}W_{t_{1}}\\ &\times\prod\limits_{i=2}^{k}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})\rho(x_{i}){\mathrm{d}}x_{i}{\mathrm{d}}W_{t_{i}},\end{split}$
@@
(8.3) $\begin{split}u^{\delta}[f]=&\int_{\mathbbm{R}}f(x){\mathrm{d}}x+\delta\sum\limits_{k=1}^{\infty}\beta_{\delta}^{k}\idotsint\limits_{0<t_{1}<\cdots<t_{k}<\delta^{-2}\atop x_{0},\cdots,x_{k}\in{\mathbbm{R}}}f(\delta x_{0}){\mathrm{d}}x_{0}\prod\limits_{i=1}^{k}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})\rho(x_{i}){\mathrm{d}}x_{i}{\mathrm{d}}W_{t_{i}},\end{split}$
@@
(8.4) $\begin{split}{\mathbbm{E}}\big{[}(u^{\delta}[f])^{2}\big{]}=&\Big{(}\int_{\mathbbm{R}}f(x){\mathrm{d}}x\Big{)}^{2}+\delta^{2}\sum\limits_{k=1}^{\infty}\beta_{\delta}^{2k}\idotsint\limits_{0<t_{1}<\cdots<t_{k}<\delta^{-2}\atop x_{0},\cdot,x_{k},y_{0},\cdots,y_{k}\in{\mathbbm{R}}}f(\delta x_{0})f(\delta y_{0}){\mathrm{d}}x_{0}{\mathrm{d}}y_{0}\\ &\times\prod\limits_{i=1}^{k}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})g_{t_{i}-t_{i-1}}(y_{i}-y_{i-1})\rho(x_{i})\rho(y_{i}){\mathrm{d}}x_{i}{\mathrm{d}}y_{i}{\mathrm{d}}t_{i}.\end{split}$
@@
We first cite some known facts about $\widetilde{r}(t)$ from [CSZ19b, Section 8.1], which will be intensively used in the following computations. Let $r(t):[0,+\infty)\to(0,+\infty)$ such that
(8.5) $r(t)=\frac{1}{2\pi t}(1+o(1)),\quad\text{as}~{}t\to+\infty.$
@@
(8.6) ${\mathbf{P}}({\mathcal{T}}_{i}^{\delta}\in{\mathrm{d}}t)=\frac{r(t)}{{\mathcal{R}}_{\delta}}\mathbbm{1}_{[0,\delta^{-2}]}(t){\mathrm{d}}t,\quad\text{where}~{}{\mathcal{R}}_{\delta}=\int_{0}^{\delta^{-2}}r(t){\mathrm{d}}t$
is the normalization constant. It is not hard to see that ${\mathcal{T}}_{i}^{\delta}$ is a continuous analogue of $\iota^{(N)}_{1}$ (see (2.21)) Note that the following properties hold for generic $r(t)$, that is, the exact value of $o(1)$ and the value of $r(t)$ in a neighborhood of 00 is not important:
@@
    (8.7) ${\mathbf{P}}({\mathcal{T}}_{1}^{\delta}+\cdots+{\mathcal{T}}_{\lfloor s\log\delta^{-2}\rfloor}^{\delta}\leq\delta^{-2})\leq e^{s-c\log s},\quad\forall~{}\delta\in(0,1),\forall~{}s\in[0,+\infty).$
@@
    (8.8) $\lim\limits_{\delta\to 0}\frac{1}{\log\delta^{-2}}\sum\limits_{k=1}^{\infty}\lambda_{\delta}^{k}{\mathbf{P}}({\mathcal{T}}_{1}^{\delta}+\cdots+{\mathcal{T}}_{k}^{\delta}\leq\delta^{-2}T)=\int_{0}^{\infty}e^{\vartheta u}{\mathbf{P}}(Y_{u}\leq T){\mathrm{d}}t,$
    where $(Y_{s})_{s\geq 0}$ is the Dickman subordinator with density $\alpha_{s}(t)$ (see (2.25)).
Note that by (8.7), the main contribution in (8.8) comes from $k\leq A(\log\delta^{-1})$ for large $A$. Moreover, we can further restrict the summation to “no consecutive increments less than $(\log\delta^{-1})^{2}$, since
(8.9) ${\mathbf{P}}\Big{(}\bigcup\limits_{k=1}^{A\log\delta^{-1}}\!\!\!\!\big{\{}{\mathcal{T}}_{k}^{\delta}\leq(\log\delta^{-1})^{2},{\mathcal{T}}_{k+1}^{\delta}\leq(\log\delta^{-1})^{2}\big{\}}\Big{)}\leq A\log\delta^{-1}{\mathbf{P}}\big{(}{\mathcal{T}}_{1}^{\delta}\leq(\log\delta^{-1})^{2}\big{)}^{2}\leq C_{A}\Big{(}\frac{\log\log\delta^{-1}}{\log\delta^{-1}}\Big{)}^{2}.$
We now compute the limit for the second moment of $u^{\delta}[f]$ as $\delta\to 0$.
8.2. Second moment of $u^{\delta}[f]$
We only need to compute the limit for the summation on the right-hand side of (8.4). Our approach is an adaption of [CSZ19b, Section 8].

### v3

@@ 6.1 Moment bounds for the coarse-grained variable $\Theta_{N,\varepsilon}$
We only need to treat $k=2,3,4$. Note that since $\Theta^{\rm(G)}$ is Gaussian and has the same covariance structure as $\Theta$, we have that
${\mathcal{E}}_{G}^{1}(\vec{\mathrm{i}})=0,\quad{\mathcal{E}}_{G}^{2}(\vec{\mathrm{i}})={\mathcal{E}}_{\Theta}^{2}(\vec{\mathrm{i}}),\quad{\mathcal{E}}_{G}^{3}(\vec{\mathrm{i}})=0,\quad{\mathcal{E}}_{G}^{4}(\vec{\mathrm{i}})=3({\mathcal{E}}_{G}^{2}(\vec{\mathrm{i}}))^{2}=3({\mathcal{E}}_{\Theta}^{2}(\vec{\mathrm{i}}))^{2}.$
For $k=2,4$, by the proof of Lemma 6.1, we have that
${\mathcal{E}}_{W}^{k}(\vec{\mathrm{i}})\leq C{\mathcal{E}}_{W}^{k}((\mathrm{i},\mathrm{i}))g_{\mathrm{i}^{\prime}-\mathrm{i}}^{k}(0){\mathcal{E}}_{W}^{k}((\mathrm{i}^{\prime},\mathrm{i}^{\prime}))={\mathbbm{E}}\bigg[\bigg(W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}((\mathrm{i},\mathrm{i}))g_{\mathrm{i}^{\prime}-\mathrm{i}}(0)W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}((\mathrm{i}^{\prime},\mathrm{i}^{\prime}))\bigg)^{k}\bigg].$
For $k=3$, we have that ${\mathcal{E}}_{W}^{3}(\vec{(}\mathrm{i}))=c_{1}^{3}{\mathcal{E}}_{\Theta}^{3}(\vec{\mathrm{i}})$ and
@@
$\begin{split}{\mathbbm{E}}\bigg[\bigg(W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}((\mathrm{i},\mathrm{i}))g_{\mathrm{i}^{\prime}-\mathrm{i}}(0)W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}((\mathrm{i}^{\prime},\mathrm{i}^{\prime}))\bigg)^{3}\bigg]&=c_{1}^{6}{\mathcal{E}}_{W}^{3}((\mathrm{i},\mathrm{i}))g_{\mathrm{i}^{\prime}-\mathrm{i}}^{3}(0){\mathcal{E}}_{W}^{3}((\mathrm{i}^{\prime},\mathrm{i}^{\prime}))\\ &=c_{1}^{6}{\mathcal{E}}_{\Theta}^{3}((\mathrm{i},\mathrm{i}))g_{\mathrm{i}^{\prime}-\mathrm{i}}^{3}(0){\mathcal{E}}_{\Theta}^{3}((\mathrm{i}^{\prime},\mathrm{i}^{\prime})).\end{split}$
In the proof of Theorem 6.3, we have explained that $|{\mathcal{E}}_{\Theta}^{3}(\vec{\mathrm{i}})|\leq C|{\mathcal{E}}_{\Theta}^{3}((\mathrm{i},\mathrm{i}))g_{\mathrm{i}^{\prime}-\mathrm{i}}^{3}(0){\mathcal{E}}_{\Theta}^{3}((\mathrm{i}^{\prime},\mathrm{i}^{\prime}))|$, and thus, the decoupling procedure can also be transferred here.
@@
Finally, by the above treatment, the upper bound of (7.17) has the same structure as $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\Theta_{n})$. Proposition 6.5 also holds given that we have the same second and fourth moments bounds on $W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}})$. We are left to obtain analogues of Lemma 6.1 and Lemma 6.2.
@@
For ${\mathcal{E}}_{W}^{4}(\vec{\mathrm{i}})$, we have that by Jensen’s inequality,
${\mathcal{E}}_{W}^{4}(\vec{\mathrm{i}})=c_{1}^{4}{\mathcal{E}}_{\Theta}^{4}(\vec{\mathrm{i}})+6c_{1}^{2}c_{2}^{2}{\mathcal{E}}_{\Theta}^{2}(\vec{\mathrm{i}}){\mathcal{E}}_{G}^{2}(\vec{\mathrm{i}})+c_{2}^{4}{\mathcal{E}}_{G}^{4}(\vec{\mathrm{i}})\leq(c_{1}^{4}+6c_{1}^{2}c_{2}^{2}+3c_{2}^{4}){\mathcal{E}}_{\Theta}^{4}(\vec{\mathrm{i}}).$
Note that ${\mathcal{E}}_{\Theta}^{2}(\vec{\mathrm{i}})$ and ${\mathcal{E}}_{\Theta}^{4}(\vec{\mathrm{i}})$ have been bounded by Lemma 6.1 and Lemma 6.2.
Since $c_{1},c_{2}$ can be $su\sqrt{t},u\sqrt{t}$ or $\sqrt{t}$ for $s,u,t\in[0,1]$, all constants here are bounded and the proof is completed. ∎
8 Proof for Theorem 1.10
For simplicity, we only consider the measure $u^{\delta}({\mathrm{d}}x):=\int_{y\in{\mathbbm{R}}}u^{\delta}_{0,1}({\mathrm{d}}x,{\mathrm{d}}y)$. We write
$u^{\delta}[f]:=\iint_{{\mathbbm{R}}^{2}}f(x)u^{\delta}_{0,1}({\mathrm{d}}x,{\mathrm{d}}y).$ (8.1)
We show that $u^{\delta}[f]$ can be approximated by a coarse-grained model $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(f,\mathbbm{1}_{\mathbbm{R}})$ (similar to (2.46)) that will be specified later. Then the proof is completed by the argument in Section 7. To proceed, we will first introduce some preliminaries in the continuous setting and compute the second moment of $u^{\delta}[f]$.
8.1 Preliminaries
@@
$\begin{split}u^{\delta}[f]=&\int_{\mathbbm{R}}f(x){\mathrm{d}}x+\sum\limits_{k=1}^{\infty}\beta_{\delta}^{k}\idotsint\limits_{0<t_{1}<\cdots<t_{k}<\delta^{-2}\atop x_{0},\cdots,x_{k}\in{\mathbbm{R}}}f(x_{0})g_{t_{1}}\big(x_{1}-\frac{x_{0}}{\delta}\big)\rho(x_{1}){\mathrm{d}}x_{0}{\mathrm{d}}x_{1}{\mathrm{d}}W_{t_{1}}\\ &\times\prod\limits_{i=2}^{k}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})\rho(x_{i}){\mathrm{d}}x_{i}{\mathrm{d}}W_{t_{i}},\end{split}$ (8.2)
@@
$\begin{split}u^{\delta}[f]=&\int_{\mathbbm{R}}f(x){\mathrm{d}}x+\delta\sum\limits_{k=1}^{\infty}\beta_{\delta}^{k}\idotsint\limits_{0<t_{1}<\cdots<t_{k}<\delta^{-2}\atop x_{0},\cdots,x_{k}\in{\mathbbm{R}}}f(\delta x_{0}){\mathrm{d}}x_{0}\prod\limits_{i=1}^{k}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})\rho(x_{i}){\mathrm{d}}x_{i}{\mathrm{d}}W_{t_{i}},\end{split}$ (8.3)
@@
$\begin{split}{\mathbbm{E}}\big[(u^{\delta}[f])^{2}\big]=&\Big(\int_{\mathbbm{R}}f(x){\mathrm{d}}x\Big)^{2}+\delta^{2}\sum\limits_{k=1}^{\infty}\beta_{\delta}^{2k}\idotsint\limits_{0<t_{1}<\cdots<t_{k}<\delta^{-2}\atop x_{0},\cdot,x_{k},y_{0},\cdots,y_{k}\in{\mathbbm{R}}}f(\delta x_{0})f(\delta y_{0}){\mathrm{d}}x_{0}{\mathrm{d}}y_{0}\\ &\times\prod\limits_{i=1}^{k}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})g_{t_{i}-t_{i-1}}(y_{i}-y_{i-1})\rho(x_{i})\rho(y_{i}){\mathrm{d}}x_{i}{\mathrm{d}}y_{i}{\mathrm{d}}t_{i}.\end{split}$ (8.4)
@@
We first cite some known facts about $\widetilde{r}(t)$ from [14, Section 8.1], which will be intensively used in the following computations. Let $r(t):[0,+\infty)\to(0,+\infty)$ such that
$r(t)=\frac{1}{2\pi t}(1+o(1)),\quad\text{as}~t\to+\infty.$ (8.5)
@@
${\mathbf{P}}({\mathcal{T}}_{i}^{\delta}\in{\mathrm{d}}t)=\frac{r(t)}{{\mathcal{R}}_{\delta}}\mathbbm{1}_{[0,\delta^{-2}]}(t){\mathrm{d}}t,\quad\text{where}~{\mathcal{R}}_{\delta}=\int_{0}^{\delta^{-2}}r(t){\mathrm{d}}t$ (8.6)
is the normalization constant. It is not hard to see that ${\mathcal{T}}_{i}^{\delta}$ is a continuous analogue of $\iota^{(N)}_{1}$ (see (2.21)). Note that the following properties hold for generic $r(t)$, that is, the exact value of $o(1)$ and the value of $r(t)$ in a neighborhood of 0 is not important:
@@
    ${\mathbf{P}}({\mathcal{T}}_{1}^{\delta}+\cdots+{\mathcal{T}}_{\lfloor s\log\delta^{-2}\rfloor}^{\delta}\leq\delta^{-2})\leq e^{s-c\log s},\quad\forall~\delta\in(0,1),\forall~s\in[0,+\infty).$ (8.7)
@@
    $\lim\limits_{\delta\to 0}\frac{1}{\log\delta^{-2}}\sum\limits_{k=1}^{\infty}\lambda_{\delta}^{k}{\mathbf{P}}({\mathcal{T}}_{1}^{\delta}+\cdots+{\mathcal{T}}_{k}^{\delta}\leq\delta^{-2}T)=\int_{0}^{\infty}e^{\vartheta u}{\mathbf{P}}(Y_{u}\leq T){\mathrm{d}}u,$ (8.8)
    where $(Y_{s})_{s\geq 0}$ is the Dickman subordinator with density $f_{s}(t)$ (see (2.25)).
Note that by (8.7), the main contribution in (8.8) comes from $k\leq A(\log\delta^{-1})$ for large $A$. Moreover, we can further restrict the summation to “no consecutive increments less than $(\log\delta^{-1})^{2}$”, since by a union bound and (8.6),
$\begin{split}&{\mathbf{P}}\Big(\bigcup\limits_{k=1}^{A\log\delta^{-1}}\big\{{\mathcal{T}}_{k}^{\delta}\leq(\log\delta^{-1})^{2},{\mathcal{T}}_{k+1}^{\delta}\leq(\log\delta^{-1})^{2}\big\}\Big)\\ \leq&A\log\delta^{-1}{\mathbf{P}}\big({\mathcal{T}}_{1}^{\delta}\leq(\log\delta^{-1})^{2}\big)^{2}\leq C_{A}\frac{(\log\log\delta^{-1})^{2}}{\log\delta^{-1}}.\end{split}$ (8.9)
We now compute the limit of the second moment of $u^{\delta}[f]$ as $\delta\to 0$.
8.2 Second moment of $u^{\delta}[f]$
We only need to compute the limit for the summation on the right-hand side of (8.4). Our approach is an adaption of [14, Section 8].